# Analysis of Semi-Supervised Learning on Hypergraphs

### v1

@@ 4.2.2 Pointwise convergence
where we used the change of variables $z_{j}=(x_{j}-x_{i_{0}})/\varepsilon_{n}$ for $1\leq j\leq k$ for (20), Lemmas 4.5 and 4.6 for (21), Assumption W.1 as well as the fact that $f(z_{1},\dots,z_{k}):=\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k}$ is odd symmetric for (22), and recalling $\psi(t)=|t|^{p-2}t$. Let $O$ be the orthogonal matrix so that $Oe_{d}=\nabla u(x_{i_{0}})/\|\nabla u(x_{i_{0}})\|_{2}$ where $e_{d}=(0,\dots,0,1)\in\mathbb{R}^{d}$. By the change of variables $\tilde{z}_{j}=O^{\top}z_{j}$ for $1\leq j\leq k$ and noting that $z_{1}\cdot\nabla u(x_{i_{0}})=\tilde{z}_{1}\cdot O^{\top}\nabla u(x_{i_{0}})=(\tilde{z}_{1})_{d}\|\nabla u(x_{i_{0}})\|_{2}$, we can continue our computation from (23):
@@
As in [14], by taking a union bound on all $x_{i_{0}}\in\Omega_{n}\cap\Omega^{\prime}$, we obtain that
$\biggl|\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)(x_{i_{0}})-\Delta_{\infty}^{(k,p)}(u)(x_{i_{0}})\biggr|\leq\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
@@ 4.2.3 $\Gamma$-convergence
$\Gamma$-convergence of the nonlocal energies
@@
We first assume that $\rho$ is Lipschitz. We will be in the same setting as in [37, Theorem 4.1] and therefore let $\Omega^{\prime}$ be compactly contained in $\Omega$. This implies that there exists $\delta^{\prime}>0$ such that $\Omega^{\prime\prime}:=\bigcup_{x\in\Omega^{\prime}}B(x,\delta^{\prime})\subset\Omega$. Furthermore, let $J$ be a positive mollifier supported in $\overline{B(0,1)}$ and for $0<\delta<\delta^{\prime}$ as well as $v\in\mathrm{L}^{p}(\mu)$ we set
@@
By [57, Theorem C.16], we have that $v_{\delta}\to v$ in $\mathrm{L}^{p}(\mu)$, $v_{\delta}$ are smooth and in particular, by Young’s convolution inequality, for $\ell\in\{1,2\}$,
@@
Dealing with the case where $\rho$ is not Lipschitz is done analogously to the proof of [37, Theorem 4.1], i.e. relying on the approximation of continuous functions by a monotone sequence of Lipschitz functions and the monotone convergence theorem to deduce the result. ∎
@@
We start by noting that (60) is trivial if $\mathcal{E}_{\infty}^{(k,p)}(u)=\infty$ so that we assume $u\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (60) on a dense subset of $\mathrm{W}^{1,p}(\Omega)$, namely $\mathrm{C}^{\infty}_{\mathrm{c}}(\Omega)$.
@@
In order to consider general $\rho$, we proceed as in [37, Theorem 4.1] which concludes the proof. ∎
$\Gamma$-convergence of the discrete energies
@@
We follow the three-step decomposition of [37, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}-\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [80, Lemma 4.2], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and
@@
Our next step is to assume that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above. Then, as in [37, Theorem 1.1], we use the linearity of the integral to obtain (64).
Our final step is to let $\eta$ be a general function satisfying Assumption W.1. Then, as in [37, Theorem 1.1], we use the monotone convergence theorem and approximation of $\eta$ by functions as in the previous step to obtain (65).
@@
We start by noting that (66) is trivial if $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)=\infty$ so that we assume $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)<\infty$ which implies that $\nu=\mu$, $v\in\mathrm{W}^{1,p}(\Omega)$ and $v(x_{i})=\ell_{i}$ for all $i\leq N$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (66) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=\ell_{i}$ for all $i\leq N$. We let $\nu_{n}=\mu_{n}$ and $v_{n}=v|_{\Omega_{n}}$ so $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$ and (66) is equivalent to
@@
The fact that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$ follows analogously from what is shown in [90, Proposition 4.17]: it relies on the fact that $v$ is uniformly continuous as well as $\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}\to 0$.
We follow the three-step decomposition of [37, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}+\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [37, Theorem 1.1], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and similarly to the previous proposition
@@
Next we proceed as in Proposition 4.12 or [37, Theorem 1.1]: by assuming that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above, we use the linearity of the integral to deduce (66); assuming that $\eta$ is a general function satisfying Assumption W.1, we approximate $\eta$ by functions of the type considered in the previous step and use the monotone convergence theorem to conclude. ∎
@@
The following proofs use arguments from [80]. For Proposition 4.15, the sum of all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ has to be considered directly as we plan to use the uniform convergence results for the $k=1$ case from [80, Lemma 4.5]. In Proposition 4.16, we show that $n\varepsilon_{n}^{p}\to\infty$ is the common lower bound for all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ in order for them to converge to the ill-posed continuum objective functions.
@@
With probability one, we can assume that the conclusions of Proposition 4.12 and [80, Lemma 4.5] hold.
@@
First, by the same argument as in Proposition 4.12, we can assume that $\sup_{n\geq 1}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq C$ and therefore $\nu_{n}=\mu_{n}$ and $\nu=\mu$. In particular, we also have that $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(v_{n})\leq C$ and, by [80, Lemma 4.5], we deduce the existence of a continuous function $\hat{v}$ such that for any $\Omega^{\prime}\subset\subset\Omega$, $\max_{\{i\leq n_{k}\,|\,x_{i}\in\Omega^{\prime}\}}|v_{n_{k}}(x_{i})-\hat{v}(x_{i})|\to 0$: this implies that $\hat{v}(x_{i})=\ell_{i}$ for all $i\leq N$ with probability one. We also note that $v=\hat{v}$ (in particular, $v(x_{i})=\ell_{i}$ for all $i\leq N$) and (72) reduces to proving
@@
We start by noting that (73) is trivial if $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))=\infty$ so that we assume $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))<\infty$ which implies that $\nu=\mu$ and $v\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [37, Remark 2.7], so it is sufficient to verify (73) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$. We let $\nu_{n}=\mu_{n}$ and $\hat{v}_{n}=v|_{\Omega_{n}}$.
@@
Again, we have $\|v_{n}\|_{\mathrm{L}^{\infty}}\leq C$ and, using the arguments of [90, Proposition 4.24], we can show that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$. Since $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$, in order to show (73), it therefore is sufficient to show that
@@
Compactness and proof of Theorem 3.4
@@
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line. We are going to follow the proof of [90, Lemma 4.25]
@@
With probability one, we can assume that the conclusions of Lemmas 4.18 and 4.19 hold.
@@
By Lemma 4.19, we know that there exists $C$ such that $\sup_{n>0}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(k,p)}((\mu_{n},u_{n}))<C$ and in particular, $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(u_{n})$ is uniformly bounded. Furthermore, analogously to what is described in the proof of [80, Theorem 2.1], we know that $\sup_{n>0}\|u_{n}\|_{\mathrm{L}^{\infty}}<C$ with probability 1. We can therefore apply [80, Proposition 4.4] to obtain a subsequence $\{n_{r}\}_{r=1}^{\infty}$ and $(\mu,u)\in\mathrm{TL}^{p}(\Omega)$ such that $(\mu_{n_{r}},u_{n_{r}})\to(\mu,u)$ in $\mathrm{TL}^{p}(\Omega)$.
@@
    Since $n\varepsilon^{p}_{n}\to 0$, by [80, Lemma 4.5], we know that $u$ is continuous and, for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n_{r}\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n_{r}}(x_{s})|\to 0$ and, with probability 1, $u(x_{i})=\ell_{i}$ for all $i\leq N$. By Lemma 4.18 and Proposition 2.6, we also have that $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$. Finally, by the uniqueness of the minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$, we conclude that the whole sequence $(\mu_{n},u_{n})$ converges to $(\mu,n)$ in $\mathrm{TL}^{p}(\Omega)$ and for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n}(x_{s})|\to 0$.

### v2

@@ 4.2.2 Equivalent representation of the continuum Laplacian
where we used the change of variables $z_{j}=(x_{j}-x_{i_{0}})/\varepsilon_{n}$ for $1\leq j\leq k$ for (55), Lemmas 4.3 and 4.4 for (56), Assumption W.1 as well as the fact that $f(z_{1},\dots,z_{k}):=\tilde{\eta}_{\mathrm{p}}(z_{1},\dots,z_{k})|z_{1}\cdot\nabla u(x_{i_{0}})|^{p-2}z_{1}\cdot\nabla u(x_{i_{0}})\rho(x_{i_{0}})^{k}$ is odd symmetric for (57), and recalling $\psi(t)=|t|^{p-2}t$. Let $O$ be the orthogonal matrix so that $Oe_{d}=\nabla u(x_{i_{0}})/\|\nabla u(x_{i_{0}})\|_{2}$ where $e_{d}=(0,\dots,0,1)\in\mathbb{R}^{d}$. By the change of variables $\tilde{z}_{j}=O^{\top}z_{j}$ for $1\leq j\leq k$ and noting that $z_{1}\cdot\nabla u(x_{i_{0}})=\tilde{z}_{1}\cdot O^{\top}\nabla u(x_{i_{0}})=(\tilde{z}_{1})_{d}\|\nabla u(x_{i_{0}})\|_{2}$, we can continue our computation from (58):
@@
As in [15], by taking a union bound on all $x_{i_{0}}\in\Omega_{n}\cap\Omega^{\prime}$ and using Corollary 4.9, we obtain that
$\biggl|\Delta_{n,\varepsilon_{n}}^{(k,p)}(u)(x_{i_{0}})-\rho(x_{i_{0}})\Delta_{\infty}^{(k,p)}(u)(x_{i_{0}})\biggr|\leq\mathcal{O}\left(\delta\|u\|_{\mathrm{C}^{3}(\mathbb{R}^{d})}^{p-1}\right)$
@@ 4.3 $\Gamma$-convergence
4.3.1 $\Gamma$-convergence of the nonlocal energies
@@
We first assume that $\rho$ is Lipschitz. We will be in the same setting as in [39, Theorem 4.1] and therefore let $\Omega^{\prime}$ be compactly contained in $\Omega$. This implies that there exists $\delta^{\prime}>0$ such that $\Omega^{\prime\prime}:=\bigcup_{x\in\Omega^{\prime}}B(x,\delta^{\prime})\subset\Omega$. Furthermore, let $J$ be a positive mollifier supported in $\overline{B(0,1)}$ and for $0<\delta<\delta^{\prime}$ as well as $v\in\mathrm{L}^{p}(\mu)$ we set
@@
By [61, Theorem C.16], we have that $v_{\delta}\to v$ in $\mathrm{L}^{p}(\mu)$, $v_{\delta}$ are smooth and in particular, by Young’s convolution inequality, for $\ell\in\{1,2\}$,
@@
Dealing with the case where $\rho$ is not Lipschitz is done analogously to the proof of [39, Theorem 4.1], i.e. relying on the approximation of continuous functions by a monotone sequence of Lipschitz functions and the monotone convergence theorem to deduce the result. ∎
@@
We start by noting that (95) is trivial if $\mathcal{E}_{\infty}^{(k,p)}(u)=\infty$ so that we assume $u\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [39, Remark 2.7], so it is sufficient to verify (95) on a dense subset of $\mathrm{W}^{1,p}(\Omega)$, namely $\mathrm{C}^{\infty}_{\mathrm{c}}(\Omega)$.
@@
In order to consider general $\rho$, we proceed as in [39, Theorem 4.1] which concludes the proof. ∎
4.3.2 $\Gamma$-convergence of the discrete energies
@@
We follow the three-step decomposition of [39, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}-\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [84, Lemma 4.2], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and
@@
Our next step is to assume that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above. Then, as in [39, Theorem 1.1], we use the linearity of the integral to obtain (99).
Our final step is to let $\eta$ be a general function satisfying Assumption W.1. Then, as in [39, Theorem 1.1], we use the monotone convergence theorem and approximation of $\eta$ by functions as in the previous step to obtain (100).
@@
We start by noting that (101) is trivial if $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)=\infty$ so that we assume $\mathcal{F}_{\infty}^{(k,p)}(\nu,v)<\infty$ which implies that $\nu=\mu$, $v\in\mathrm{W}^{1,p}(\Omega)$ and $v(x_{i})=\ell_{i}$ for all $i\leq N$. Furthermore, we are going to apply [39, Remark 2.7], so it is sufficient to verify (101) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=\ell_{i}$ for all $i\leq N$. We let $\nu_{n}=\mu_{n}$ and $v_{n}=v|_{\Omega_{n}}$ so $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$ and (101) is equivalent to
@@
The fact that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$ follows analogously from what is shown in [94, Proposition 4.17]: it relies on the fact that $v$ is uniformly continuous as well as $\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}\to 0$.
We follow the three-step decomposition of [39, Theorem 1.1]. First, suppose that $\eta(t)=a$ if $0\leq t\leq b$ and $\eta(t)=0$ else where $a$ and $b$ are positive constants. Define $\tilde{\varepsilon}_{n}=\varepsilon_{n}+\frac{2\|T_{n}-\mathrm{Id}\|_{\mathrm{L}^{\infty}}}{b}$. From [39, Theorem 1.1], we know that $\frac{\varepsilon_{n}}{\tilde{\varepsilon}_{n}}\to 1$ and similarly to the previous proposition
@@
Next we proceed as in Proposition 4.15 or [39, Theorem 1.1]: by assuming that $\eta=\sum_{k=1}^{\ell}\eta_{l}$ satisfies Assumption W.1 where $\eta_{l}$ are functions of the type considered in the above, we use the linearity of the integral to deduce (101); assuming that $\eta$ is a general function satisfying Assumption W.1, we approximate $\eta$ by functions of the type considered in the previous step and use the monotone convergence theorem to conclude. ∎
@@
The following proofs use arguments from [84]. For Proposition 4.18, the sum of all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ has to be considered directly as we plan to use the uniform convergence results for the $k=1$ case from [84, Lemma 4.5]. In Proposition 4.19, we show that $n\varepsilon_{n}^{p}\to\infty$ is the common lower bound for all energies $\{\mathcal{F}_{n,\varepsilon_{n}}^{(k,p)}\}_{k=1}^{q}$ in order for them to converge to the ill-posed continuum objective functions.
@@
With probability one, we can assume that the conclusions of Proposition 4.15 and [84, Lemma 4.5] hold.
@@
First, by the same argument as in Proposition 4.15, we can assume that $\sup_{n\geq 1}\left(\mathcal{S}\mathcal{F}\right)_{n,\varepsilon_{n}}^{(k,p)}((\nu_{n},v_{n}))\leq C$ and therefore $\nu_{n}=\mu_{n}$ and $\nu=\mu$. In particular, we also have that $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(v_{n})\leq C$ and, by [84, Lemma 4.5], we deduce the existence of a continuous function $\hat{v}$ such that for any $\Omega^{\prime}\subset\subset\Omega$, $\max_{\{i\leq n_{k}\,|\,x_{i}\in\Omega^{\prime}\}}|v_{n_{k}}(x_{i})-\hat{v}(x_{i})|\to 0$: this implies that $\hat{v}(x_{i})=\ell_{i}$ for all $i\leq N$ with probability one. We also note that $v=\hat{v}$ (in particular, $v(x_{i})=\ell_{i}$ for all $i\leq N$) and (107) reduces to proving
@@
We start by noting that (108) is trivial if $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))=\infty$ so that we assume $\mathcal{G}_{\infty}^{(k,p)}((\nu,v))<\infty$ which implies that $\nu=\mu$ and $v\in\mathrm{W}^{1,p}(\Omega)$. Furthermore, we are going to apply [39, Remark 2.7], so it is sufficient to verify (108) on a dense subset of $\{\mu\}\times\mathrm{W}^{1,p}(\Omega)$, namely we consider $(\mu,v)\in\{\mu\}\times\mathrm{C}^{\infty}(\bar{\Omega})$. We let $\nu_{n}=\mu_{n}$ and $\hat{v}_{n}=v|_{\Omega_{n}}$.
@@
Again, we have $\|v_{n}\|_{\mathrm{L}^{\infty}}\leq C$ and, using the arguments of [94, Proposition 4.24], we can show that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{p}(\Omega)$. Since $v_{n}(x_{i})=\ell_{i}$ for all $i\leq N$, in order to show (108), it therefore is sufficient to show that
@@
4.3.3 Compactness and proof of Theorem 3.3
@@
In the proof $C>0$ will denote a constant that can be arbitrarily large, is independent of $n$, and that may change from line to line. We are going to follow the proof of [94, Lemma 4.25]
@@
With probability one, we can assume that the conclusions of Lemmas 4.21 and 4.22 hold.
@@
By Lemma 4.22, we know that there exists $C$ such that $\sup_{n>0}(\mathcal{S}\mathcal{F})_{n,\varepsilon_{n}}^{(k,p)}((\mu_{n},u_{n}))<C$ and in particular, $\mathcal{E}_{n,\varepsilon_{n}}^{(1,p)}(u_{n})$ is uniformly bounded. Furthermore, analogously to what is described in the proof of [84, Theorem 2.1], we know that $\sup_{n>0}\|u_{n}\|_{\mathrm{L}^{\infty}}<C$ with probability 1. We can therefore apply [84, Proposition 4.4] to obtain a subsequence $\{n_{r}\}_{r=1}^{\infty}$ and $(\mu,u)\in\mathrm{TL}^{p}(\Omega)$ such that $(\mu_{n_{r}},u_{n_{r}})\to(\mu,u)$ in $\mathrm{TL}^{p}(\Omega)$.
@@
    Since $n\varepsilon^{p}_{n}\to 0$, by [84, Lemma 4.5], we know that $u$ is continuous and, for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n_{r}\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n_{r}}(x_{s})|\to 0$ and, with probability 1, $u(x_{i})=\ell_{i}$ for all $i\leq N$. By Lemma 4.21 and Proposition 2.6, we also have that $(\mu,u)$ is a minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$. Finally, by the uniqueness of the minimizer of $(\mathcal{S}\mathcal{F})_{\infty}^{(q,p)}$, we conclude that the whole sequence $(\mu_{n},u_{n})$ converges to $(\mu,n)$ in $\mathrm{TL}^{p}(\Omega)$ and for every $\Omega^{\prime}\subset\subset\Omega$, we have that $\max_{\{s\leq n\,|\,x_{s}\in\Omega^{\prime}\}}|u(x_{s})-u_{n}(x_{s})|\to 0$.